7x – 4y = –7
5x + y = 22

Answer by macston(5194)

.
What exactly is your question?
.
5x+y=22
y=22-5x
.
7x–4y=–7
7x-4(22-5x)=-7
7x-88+20x=-7
27x=81
x=3
.
5x+y=22
5(3)+y=22
15+y=22
y=7
.
Possible answer: (3,7)